The Watson Foundation, a private not-for-profit entity, starts 2017 with cash of $113,000; contributions receivable (net) of $213,000; investments of $313,000; and land, buildings, and equipment of $213,000. In addition, its unrestricted net assets were $426,000, temporarily restricted net assets were $113,000, and permanently restricted net assets were $313,000. Of the temporarily restricted net assets, 50 percent must be used to help pay for a new building; the remainder is restricted for salaries. No implied time restriction was designated for the building when purchased. For the permanently restricted net assets, all income is unrestricted.

During the current year, the entity has the following transactions:

  • Computed interest of $33,000 on the contributions receivable.
  • Received cash of $113,000 on the contributions and wrote off another $5,300 as uncollectible.
  • Received unrestricted cash gifts of $193,000.
  • Paid salaries of $103,000 with $28,000 of that amount coming from restricted funds.
  • Received a cash gift of $25,000 that the entity must convey to another charity. However, Watson has the right to give the money to a different organization if it so chooses.
  • Bought a building for $513,000 by signing a long-term note for $456,500 and using restricted funds for the remainder.
  • Collected membership dues of $43,000. Individuals receive substantial benefits from the memberships.
  • Received income of $43,000 generated by the permanently restricted net assets.
  • Paid rent of $25,000, advertising of $28,000, and utilities of $29,000.
  • Received an unrestricted pledge of $213,000; it will be collected in five years. The organization expects to collect the entire amount. Present value is $162,000. It then recognized interest of $7,300 for the year.
  • Computed depreciation as $53,000.
  • Paid $28,000 in interest on the note signed to acquire the building.
  1. Prepare a statement of activities for this entity for this year.

  2. Prepare a statement of financial position for this entity at the end of this year.
Answer #1

Answer a

Particulars Unrestricted Net Assets Temporarily Restricted Net Assets Permanently Restricted Net Assets
Support
Contributions 193000 187000 (162000+25000)
Revenue
Membership dues 43000
Investment income 43000
Interest income 33,000 7,300
Total Support and Revenue 312,000 194,300
Net assets released from restriction 84,500 (28000+(513000-456500) (84,500)
Total revenues and net assets released from restriction 396,500 109,800
Expenses
General and administrative
Rent 25,000
Salary 103,000
Advertising 28,000
utilities 29,000
Depreciation 53,000
Interest 28,000
Total expenses 266,000
Excess of total revenues and net assets released from restriction over expenses 130500 109,800
Net assets at beginning of year 426000 113000 313000
Net assets at end of year 556500 222800 313000

STATEMENT OF FINANCIAL POSITION

Amount $ Total Amount $
ASSETS
Cash 260,500
Pledges receivable (net) 302,300
Investments 313000
Land, buildings, and equipment 726000
Depreciation 53000 673000
Total assets 1,548,800
LIABILITIES
Notes Payable 456500
NET ASSETS
Unrestricted 556,500
Temporarily Restricted 222,800
Permanently Restricted 313,000 1,092,300
TOTAL LIABILITIES AND NET ASSETS 1,548,800


Cash = 113000+113000+193000-103000+25000-(513000-456500)+43000+43000-25000-28000-29000-28000 = 260500

Pledges receivable = 213000+33000+162000+7300 = 415,300-113000

Land and buildings = 213000+513000